# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ Appendix F Experimental Setting of Visualization of the Attack Trajectory
We visualize the attack by plotting the loss landscape and trace the trajectories of EoT attack under White-box setting and the Deterministic White-box setting in Figure 3. We run a vanilla PGD20-EoT10 attack under White-box setting and a PGD20 attack under Deterministic White-box setting. We then expand a 2D space using the final perturbations from these two attacks, draw the loss landscape, and plot the attack trajectories on it. Note that the two adversarial perturbation directions are not strictly orthogonal. To extend this 2D space, we use the Deterministic White-box attack direction and the orthogonal component of the EoT attack direction. Note that the endpoints of both trajectories lie exactly on the loss landscape, while intermediate points are projected onto it. The plot is evaluated using WideResNet-28-10 with DPDDPM over the first 128 images of CIFAR-10 dataset.
@@ Appendix G Pseudo-code of ADDT
9:        $\bm{\delta}=\bm{\delta}+\nabla_{\bm{\epsilon}^{\prime}}L(C(\bm{P}(\bm{x}_{t},t),y))$
@@
11:     $\bm{\epsilon}\sim\mathcal{N}(0,I)$
@@
14:     Take a gradient descent step on:$\nabla_{\bm{\theta}}\|\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}}(\bm{x}_{0}-\bm{P}(\bm{x}_{t},t))\|_{2}^{2}$
15:  end forDiffusion model $\bm{\epsilon}_{\bm{\theta}}$ predicts the Gaussian noise added to the image, adopting Equation 4 in the paper, we have $\bm{P}(\bm{x}_{t},t)=\left(\bm{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\right)/\sqrt{\overline{\alpha}_{t}}$
@@ Appendix H ADDT Results on DPDDIM
As shown in Table 12, the performance of DPDDIM is less sensitive to the number of function evaluations (NFE). Additionally, ADDT consistently improved the robustness of DPDDIM.
Table 12: Clean and robust accuracy (%) on DPDDIM. ADDT improve robustness across different NFE (*: default DDIM generation setting, -: classifier only ).
Dataset NFE Vanilla ADDT
$Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$
@@ Appendix I Adopting VPSDE(DDPM++) and EDM Models in DBP
From a unified perspective, diffusion processes can be modeled by stochastic differential equations (SDE) (Song et al., 2021). The forward SDE, as described in Equation 14, converts a complex initial data distribution into a simpler, predetermined prior distribution by progressively infusing noise. This can also be done in a single step, as shown in Equation 15, mirroring the strategy of DDPM described in Equation 2. Reverse SDE, as explained in Equation 16, reverses this process, restoring the noise distribution to the original data distribution, thus completing the diffusion cycle.
@@
The reverse process of SDEs also derives equivalent ODEs Equation 17 for fast sampling and exact likelihood computation, and this Score ODEs corresponds to DDIM.
@@
EDM provides a unified framework to synthesize the design principles of different diffusion models (DDPM,DDIM,iDDPM (Nichol & Dhariwal, 2021),VPSDE,VESDE (Song et al., 2021)). Within this framework, EDM incorporates efficient sampling methods, such as the Heun sampler, and introduces optimized scheduling functions $\sigma(t)$ and $s(t)$. This allows EDM to achieve state-of-the-art performance in generative tasks.
@@
$\bm{x}_{t}=\bm{x}_{0}+\sigma(t^{*})*\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),$ (18)
where we choose $\sigma(t^{*})=0.5$ for clean and robust accuracy tradeoff. And for reverse process, EDM incorporates a parameter $S_{churn}$ to modulate the stochastic noise infused during the reverse process. For our experiments, we choose 50 reverse steps (50 NFE, NFE is Function of Neural Function Evaluations), configured the parameters with $S_{min}=0.01$, $S_{max}=0.46$, $S_{noise}=1.007$, and designate $S_{churn}=0$ to represent EDM-ODE, $S_{churn}=6$ to represent EDM-SDE.
@@
Type Vanilla ADDT
@@
Song et al. present a Predictor-Corrector sampler for SDEs reverse process for VPSDE (DDPM++) (as detailed in Appendix I of Song et al. (2021)). However, standard implementations of VPSDE (DDPM++) typically use only the Predictor. Given our hypothesis that stochasticity contributes to robustness, we expect that integrating the Corrector sampler into VPSDE (DDPM++) would further enhance the robustness of DBP models. Our empirical results, as shown in Table 14, confirm that the inclusion of a Corrector to VPSDE (DDPM++) indeed improve the model’s defenses ability against adversarial attacks with $l_{\infty}$ norm constraints. This finding supports our claim that the increased stochasticity can further strengthen DBP robustness. Adding Corrector is also consistent with ADDT. Note that the robustness against $l_{2}$ norm attacks does not show a significant improvement with the integration of the Extra Corrector. A plausible explanation for this could be that the robustness under $l_{2}$ attacks is already quite strong, and the compromised performance on clean data counteracts the increase in robustness.
@@
Type Vanilla Extra Corrector ADDT ADDT+Extra Corrector
@@
$l_{\infty}$ 55.96 59.77 62.11 65.23
$l_{2}$ 75.78 74.22 76.66 76.66
Appendix K Evaluating RBGM-Mapped Perturbations
@@
$Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$
DPDDPM-1000 CIFAR-10 85.94 47.27 69.34 85.64 51.46 70.12 84.47 52.64 68.55
@@
[…]
Appendix L Evaluation with Fixed AutoAttack
AutoAttack (Croce & Hein, 2020), an ensemble of White-box and Black-box attacks, is a popular benchmark for evaluating model robustness. It is used in RobustBench (Croce et al., 2020) to evaluate over 120 models. However, Nie et al. Nie et al. (2022) found that the Rand version of AutoAttack, designed to evaluate stochastic defenses, sometimes yields higher accuracy than the Standard version, intended for deterministic methods. Our comparison of AutoAttack and PGD20-EoT10 in Table 16 also shows that the Rand version of AutoAttack gives higher accuracy than the PGD20-EoT10 attack.
We attribute this to the sample selection of AutoAttack. As an ensemble of attack methods, AutoAttack selects the final adversarial sample from either the original input or the attack results. However, the original implementation neglects stochasticity and considers a adversarial sample to be sufficiently adversarial if it gives a false result in one evaluation. To fix this, we propose a 20-iteration evaluation and selects the adversarial example with the lowest accuracy. The flawed code is in the official GitHub main branch, git version $a39220048b3c9f2cca9a4d3a54604793c68eca7e$, and specifically in lines #125, #129, #133-136, #157, #221-225, #227-228, #231 of the file autoattack/autoattack.py. We will open source our updated code and encourage future stochastic defense methods to be evaluated against the fixed code. The code now can be found at: https://anonymous.4open.science/r/auto-attack-595C/README.md.
After the fix, AutoAttack’s accuracy dropped by up to 10 points, producing similar results to our PGD20-EoT10 test results. However, using AutoAttack on DPDDPM with $S=1000$ took nearly 25 hours, five times longer than PGD20-EoT10, so we will use PGD20-EoT10 for the following test.
Method $l_{\infty}$ $l_{2}$
DiffPure 62.11 56.25 55.96 81.84 76.37 75.78
DPDDPM-1000 57.81 46.88 48.63 71.68 71.09 72.27
DPDDIM-100 50.20 40.62 44.73 77.15 70.70 71.68
Table 17: Clean and robust accuracy (%) on different DBP methods for CIFAR-10, evaluated with AutoAttackADDT (Rand version). All methods show consistent improvement when fine-tuned with ADDT.
Method Vanilla ADDT
$Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$
DiffPure 89.26 56.25 76.37 89.94 58.20 77.34
DPDDPM 85.94 46.88 71.09 85.64 48.63 72.27
DPDDIM 88.38 40.62 70.70 88.77 44.73 71.68
Appendix M Robustness under Deterministic White-box Setting
@@
We evaluate robustness of ADDT fine-tuned models under Deterministic White-box Setting. We assess robustness across varying NFE and compare the performance of vanilla models and ADDT fine-tuned models, with results showcased in Figure 9. ADDT consistently improves model performance at different NFE, particularly noticeable at lower NFE. This affirms that ADDT equips the diffusion model with the ability to directly counter adversarial perturbations.
[…]
@@
Figure 9: Revisiting Deterministic White-box Robustness. ADDT improves robustness under both White-box and Deterministic White-box setting, implying that ADDT strengthens the models’ ability to handle adversarial inputs. The dashed line on the right is the performance of the vanilla model.
Appendix N Ablation Study of $\lambda_{unit}$
@@
In Section 6.1 we choose $\lambda_{unit}$=0.03 because most of the adversarial perturbations are in this range. We also provide an ablation study here, which shows that the performance of ADDT is insensitive to $\lambda_{unit}$ and gets a consistent improvement.
Table 18: Ablation study of $\lambda_{unit}$, ADDT is insensitive to it and gets a consistent improvement on robust accuracy (%).
[…]
$l_{\infty}$ $Clean$ 21.78 36.72 45.21 46.78 47.27
0.02 24.02 40.92 48.14 48.83 48.93
0.03 30.27 43.07 48.44 50.20 51.46
0.04 31.25 44.92 50.68 51.07 50.88
$l_{2}$ $Clean$ 36.13 55.47 65.23 68.85 69.34
0.02 41.99 61.72 67.48 69.82 70.31
0.03 41.99 62.97 69.82 69.04 70.12
0.04 49.02 64.45 69.24 69.53 69.92
Appendix O Computational Cost of ADDT
@@
Optimizing DDPM and DDIM models through ADDT achieves near-optimal performance in 50 epochs, taking 12 hours on 4 NVIDIA GeForce RTX 2080 Ti GPUs. This matches the speed of traditional adversarial training and is significantly faster than the latest adversarial training techniques that use diffusion models to augment the dataset (Wang et al., 2023). Testing DPDDPM and DPDDIM is computationally costly due to the employment of EoT; it takes approximately 5 hours on four NVIDIA GeForce RTX 2080 Ti GPUs to validate 1,024 images on the CIFAR10/CIFAR100 datasets.
A notable feature of ADDT is its one-and-done training approach. After the initial training, ADDT can protect various classifiers without the need for further fine-tuning, as shown in Section 6.1. This differs from adversarial classifier training, which requires individual training for each classifier.
@@
The code was developed independently by two individuals and mutually verified, with consistent results achieved through independent training and testing. We will also make the code open-source and remain committed to advancing the field.

### v3

@@ Appendix F Experimental Setting of Visualization of the Attack Trajectory
We visualize the attack by plotting the loss landscape and tracing the trajectories of EoT attack under White-box setting and the Deterministic White-box setting in Figure 3. We run a vanilla PGD20-EoT10 attack under White-box setting and a PGD20 attack under Deterministic White-box setting. We then expand a 2D space using the final perturbations from these two attacks, draw the loss landscape, and plot the attack trajectories on it. Note that the two adversarial perturbation directions are not strictly orthogonal. To extend this 2D space, we use the Deterministic White-box attack direction and the orthogonal component of the EoT attack direction. Note that the endpoints of both trajectories lie exactly on the loss landscape, while intermediate points are projected onto it. The plot is evaluated using WideResNet-28-10 with DPDDPM over the first 128 images of CIFAR-10 dataset.
@@ Appendix G Pseudo-code of ADDT
9:        $\bm{\delta}=\bm{\delta}+\nabla_{\bm{\epsilon}^{\prime}}L(C(\bm{P}(\bm{\theta},\bm{x}_{t},t),y))$
@@
11:     $\bm{\epsilon}\sim\mathcal{N}\left(0,I\right)$
@@
14:     Take a gradient descent step on:$\nabla_{\bm{\theta}}\|\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}}(\bm{x}_{0}-\bm{P}(\bm{\theta},\bm{x}_{t},t))\|_{2}^{2}$
15:  end forDiffusion model $\bm{\epsilon}_{\bm{\theta}}$ predicts the Gaussian noise added to the image, adopting Equation 4 in the paper, we have $\bm{P}(\bm{\theta},\bm{x}_{t},t)=\left(\bm{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\right)/\sqrt{\overline{\alpha}_{t}}$
@@ Appendix H ADDT Results on DPDDIM
As shown in Table 14, the performance of DPDDIM is less sensitive to the number of function evaluations (NFEs). Additionally, ADDT consistently improved the robustness of DPDDIM.
Table 14: Clean and robust accuracy (%) on DPDDIM. ADDT improve robustness across different NFEs (*: default DDIM generation setting, -: classifier only ).
Dataset NFEs Vanilla ADDT
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$
@@ Appendix I Adopting VPSDE(DDPM++) and EDM Models in DBP
From a unified perspective, the diffusion process can be modeled by stochastic differential equations (SDE) (Song et al., 2021b). The forward SDE, as described in Equation 14, converts a complex initial data distribution into a simpler, predetermined prior distribution by progressively infusing noise. This can also be done in a single step, as shown in Equation 15, mirroring the strategy of DDPM described in Equation 2. Reverse SDE, as explained in Equation 16, reverses this process, restoring the noise distribution to the original data distribution, thus completing the diffusion cycle.
@@
The reverse process of SDEs also derives equivalent ODEs Equation 17 for fast sampling and exact likelihood computation, and these Score ODEs correspond to DDIM.
@@
EDM provides a unified framework to synthesize the design principles of different diffusion models (DDPM, DDIM,iDDPM (Nichol & Dhariwal, 2021), VPSDE, VESDE (Song et al., 2021b)). Within this framework, EDM incorporates efficient sampling methods, such as the Heun sampler, and introduces optimized scheduling functions $\sigma(t)$ and $s(t)$. This allows EDM to achieve state-of-the-art performance in generative tasks.
@@
$\bm{x}_{t}=\bm{x}_{0}+\sigma(t^{*})\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (18)
where we choose $\sigma(t^{*})=0.5$ for clean and robust accuracy tradeoff. And for the reverse process, EDM incorporates a parameter $S_{churn}$ to modulate the stochastic noise infused during the reverse process. For our experiments, we choose 50 reverse steps (50 NFEs, NFEs is Function of Neural Function Evaluations), configured the parameters with $S_{min}=0.01$, $S_{max}=0.46$, $S_{noise}=1.007$, and designate $S_{churn}=0$ to represent EDM-ODE, $S_{churn}=6$ to represent EDM-SDE.
@@
Metric Vanilla ADDT
@@
(Song et al., 2021b) present a Predictor-Corrector sampler for SDEs reverse process for VPSDE (DDPM++), as detailed in Appendix I. However, standard implementations of VPSDE (DDPM++) typically use only the Predictor. Given our hypothesis that stochasticity contributes to robustness, we expect that integrating the Corrector sampler into VPSDE (DDPM++) would further enhance the robustness of DBP models. Our empirical results, as shown in Table 16, confirm that the inclusion of a Corrector to VPSDE (DDPM++) indeed improves the model’s defense ability against adversarial attacks with $\ell_{\infty}$ norm constraints. This finding supports our claim that the increased stochasticity can further strengthen DBP robustness. Adding Corrector is also consistent with ADDT. Note that the robustness against $\ell_{2}$ norm attacks does not show a significant improvement with the integration of the Extra Corrector. A plausible explanation for this could be that the robustness under $\ell_{2}$ attacks is already quite strong, and the compromised performance on clean data counteracts the increase in robustness.
@@
Metric Vanilla Extra Corrector ADDT ADDT+Extra Corrector
@@
$\ell_{\infty}$ 55.96 59.77 62.11 65.23
$\ell_{2}$ 75.78 74.22 76.66 76.66
Appendix K Discussion about RBGM-Mapped Perturbations
K.1 Motivation and Advantages of RBGM
In Section 4, we discuss the limitations of Diffusion-Based Perturbation (DBP) models in effectively purifying adversarial perturbations. To overcome these limitations and simultaneously preserve the generative ability of the diffusion models, we introduce a novel approach: incorporating “adversarially selected Gaussian noise” into the diffusion training process.
To elaborate, a conventional diffusion forward process is based on the equation:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\,\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\,\bm{\epsilon},$ (19)
where $\bm{x}_{t}$ represents the noisy image at time $t$, $\bm{x}_{0}$ is the initial input, $\overline{\alpha}_{t}$ is a time-dependent scaling factor, and $\bm{\epsilon}$ is random Gaussian noise. Our proposed method, ADDT, modifies this equation to include an adversarial component:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\,\bm{x}_{0}+\sqrt{1-\lambda_{t}^{2}}\,\sqrt{1-\overline{\alpha}_{t}}\,\bm{\epsilon}+\lambda_{t}\,\sqrt{1-\overline{\alpha}_{t}}\,\bm{\epsilon}_{\bm{\delta}}(\bm{\delta}).$ (20)
In this revised formulation, $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$ represents the adversarial perturbation, and $\lambda_{t}$ is a parameter that controls the blend between traditional and adversarial noise. The core objective of ADDT training is to generate perturbations that emulate the characteristics of Gaussian noise in conventional diffusion training while incorporating adversarial disturbances. This leads to our Rank-Based Gaussian Mapping (RBGM) technique, which preserves the relative ordering of perturbation magnitudes while adjusting their values to more closely resemble a Gaussian distribution. The advantages of RBGM are twofold:
Enhancing statistical consistency.
Raw adversarial perturbation values often exhibit non-standard distributions, and RBGM serves to recalibrate these perturbations, aligning them more closely with a Gaussian distribution. To elaborate, rather than enforcing a multivariate Gaussian distribution for the entire perturbation, RBGM ensures that the distribution of individual perturbation values adheres to Gaussian characteristics.
The benefit of this transformation can be illustrated in Figure 9 and Figure 10. For a fair comparison, the perturbation values have been normalized. In Figure 9, the original perturbation values display a wide array of distributions across different images and time steps. After the mapping of RBGM, these values are transformed to exhibit a uniform Gaussian distribution.
Figure 9: Raw perturbation values exhibit diverse distributions across images and time steps. RBGM maps these perturbations to a uniform Gaussian distribution.
In Figure 10, the raw perturbations show irregular and inconsistent behavior when mixed with Gaussian noise at varying ratios. However, after RBGM mapping, the perturbations and the mixture exhibit consistent statistics with the pure Gaussian noise. The statistical consistency of the perturbation values may ease the training of the diffusion model and avoid significant deviation from the normal diffusion process.
Figure 10: RBGM ensures that mixing perturbations with Gaussian noise at any ratio yields a consistent value distribution.
Reducing image-specific dependence.
In the training of diffusion models, the Gaussian noise is independent of specific images or time steps. This characteristic contrasts with adversarial perturbations, which are typically tailored to each input. RBGM addresses this issue by introducing stochasticity into the construction of perturbations while preserving the ranks of the values of the image-dependent adversarial perturbations. This approach effectively reduces image-specific dependence. Consequently, RBGM enhances ADDT’s ability to better mimic the diffusion training process and potentially mitigates the overfitting of training images.
K.2 RBGM-Mapped Perturbations Preserve Adversarial Characteristics
While RBGM-mapped perturbations are “selected from a Gaussian distribution”, their actual distribution deviates from a pure Gaussian distribution and is adversarial for models. To substantiate this claim, we compare the influence of RBGM-mapped perturbations and Gaussian noise on model performance. In our experiments, we perturb clean images by adding RBGM-mapped perturbations and Gaussian noise, each scaled by a factor of $0.03$. The results present in Table 17 demonstrate that RBGM-mapped perturbations effectively act as adversarial inputs to the model. These perturbations drastically reduce the accuracy of a pre-trained clean WRN-28-10 from 95.12% to 4.47%.
Table 17: Comparison of model accuracy (%) under different conditions. RBGM-mapped perturbations lead to a significant reduction in accuracy compared to Gaussian noise.
Classifier Clean Gaussian noise RBGM-mapped perturbation
WRN-28-10 95.12 81.54 4.47
K.3 Blending Adversarial Perturbations into Diffusion Model Training
In conventional adversarial attacks, perturbations are directly applied to the image, resulting in an adversarial image:
$\bm{x}_{\text{adv}}=\bm{x}_{0}+\bm{\delta},$ (21)
where $\bm{x}_{0}$ is the original input image, and $\bm{\delta}$ is the adversarial perturbation. In ADDT, we incorporate this concept into the diffusion process, redefining the noisy image at time step $t$ as:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}(\bm{x}_{0}+\bm{\delta})+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},$ (22)
To enable the diffusion model to effectively purify adversarial perturbations during training, we reformulate the above equation by merging the perturbation $\bm{\delta}$ with the noise $\bm{\epsilon}$. This results in:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\,\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\,\left(\bm{\epsilon}+\gamma_{t}\,\bm{\delta}\right),$ (23)
where $\gamma_{t}$ is a scaling factor defined as:
$\gamma_{t}=\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}}.$ (24)
Since $\overline{\alpha}_{t}$ is a time-dependent parameter that monotonically decreases from $1$ to 00 during the diffusion process, $\gamma_{t}$ spans the range from 00 to $\infty$. To ensure the adversarial perturbation remains within a manageable intensity, we constrain its value to the range between $\lambda_{\text{min}}$ and $\lambda_{\text{max}}$.
K.4 RBGM Enhances Perturbation Compatibility with Diffusion Model Training
To illustrate how RBGM enhances the compatibility of perturbations with diffusion model training, we conduct comparative analyses in two scenarios. First, we assess the impact of RBGM on statistical consistency by comparing Gaussian noise with adversarial perturbations reordered based on Gaussian noise ranks. Second, we evaluate the effectiveness of RBGM-mapped perturbations in improving model robustness while maintaining generative performance by comparing them with $\ell_{2}$-normalized perturbations.
Gaussian noise vs. adversarial perturbations ordered by Gaussian noise
We begin by examining RBGM’s influence on statistical consistency through two training methodologies:
  1. 1.
    Vanilla: Trained with standard Gaussian noise.
  2. 2.
    ADDTGaussian reordered: Trained with adversarial perturbations reordered according to Gaussian noise ranks. To ensure a fair comparison, the perturbations are normalized to have a mean of 0 and a variance of 1, as their original magnitudes (derived from accumulated gradients) are significantly smaller than those of standard Gaussian noise. Note that this approach—reordering adversarial perturbations based on Gaussian noise ranks—is distinct from RBGM, where Gaussian noise is reordered based on adversarial perturbation ranks.
The results presented in Table 18 reveal that models trained with Gaussian noise reordering using adversarial perturbation values exhibit lower accuracy on both clean and adversarial samples compared to vanilla models. This decline in performance underscores RBGM’s ability to enhance perturbation compatibility with diffusion model training by improving statistical consistency.
Table 18: Comparison of DPDDPM accuracy under different conditions and perturbation types. Training with perturbations reordered by Gaussian noise and adversarial perturbation values degrades performance.
NFEs Vanilla ADDTGaussian reordered
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$
5 49.51 21.78 36.13 48.40 21.10 33.40
10 73.34 36.72 55.47 71.78 34.07 52.98
20 81.45 45.21 65.23 79.99 42.43 64.21
50 85.54 46.78 68.85 83.90 46.73 68.17
100* 85.94 47.27 69.34 84.54 46.98 69.33
RBGM-mapped perturbations vs. $\ell_{2}$-normalized perturbations
To further investigate RBGM’s effectiveness in making adversarial perturbations compatible with diffusion model training, we compare:
  1. 1.
    ADDT: Trained with RBGM-mapped perturbations.
  2. 2.
@@
As shown in Table 19, models trained with $\ell_{2}$-normalized perturbations tend to perform better under $\ell_{2}$ attacks in some scenarios (possibly because these perturbations are more similar to those generated by $\ell_{2}$ attack during testing). ADDT generally achieves better results. This advantage is particularly pronounced under $\ell_{\infty}$ attacks and in scenarios with higher NFEs. Furthermore, as shown in Table 20, ADDT yields a lower FID value, reflecting better preservation of generative capabilities.
Table 19: Comparison of DPDDPM accuracy under different perturbation conditions. Training with ADDT generally achieves better results. This advantage is particularly pronounced under $\ell_{\infty}$ attacks and in scenarios with higher NFEs.
@@
Clean $\ell_{\infty}$ $\ell_{2}$ Clean $\ell_{\infty}$ $\ell_{2}$
5 59.96 30.27 41.99 60.40 28.47 44.58
10 78.91 43.07 62.97 79.29 41.90 63.72
20 83.89 48.44 69.82 83.59 47.85 67.68
50 85.45 50.20 69.04 84.83 49.12 69.24
100* 85.64 51.46 70.12 84.97 49.95 69.29
Table 20: FID score of DDPM for CIFAR-10 fine-tuned under different conditions. Fine-tuning with ADDT results in a lower FID score compared to $\ell_{2}$ normalization.
@@
FID 3.50 5.190 5.678
As discussed in Section K.1, the goal of utilizing RGBM to generate perturbations is to mimic the characteristic of Gaussian noise, hence aligning ADDT closer to traditional diffusion model training. In this context, both RBGM and $\ell_{2}$ normalization serve as approximations of Gaussian noise. Yet, RBGM provides a more precise approximation, enhancing robustness and maintaining the generative performance more effectively than $\ell_{2}$ normalization.
Appendix L Comparing RBGM-Mapped Perturbations with $\ell_{\infty}$ Perturbations
In Section 6.3, we briefly explore the generation capabilities of diffusion models trained with RBGM-mapped and $\ell_{\infty}$ perturbations. Here, we provide further experiments and delve deeper into their robustness comparison. To train with $\ell_{\infty}$ perturbations, we adjust ADDT, replacing RBGM-mapped perturbations with $\ell_{\infty}$ perturbations. Here, instead of converting accumulated gradients to Gaussian-like perturbations, we use a 5-step projected gradient descent (PGD-5) approach. For fair comparison, we also set $\lambda_{unit}=1,\lambda_{min}=0,\lambda_{max}=10$ and refer to this modified training protocol as $\ell_{\infty}$.
We evaluate the clean and robust accuracy of ADDT and $\ell_{\infty}$ fine-tuned models. These models exhibit different behaviors. As shown in Table 21, while Gaussian-mapped perturbations can simultaneously improve clean accuracy and robustness against both $\ell_{2}$ and $\ell_{\infty}$ attacks, training with $\ell_{\infty}$ perturbations primarily improves performance against $\ell_{\infty}$ attacks.
Table 21: Clean and robust accuracy (%) on DBP models trained with different perturbations for CIFAR-10. While ADDT simultaneously improves clean accuracy and robustness against both $\ell_{2}$ and $\ell_{\infty}$ attacks. $\ell_{\infty}$ primarily improves performance against $\ell_{\infty}$ attacks.
DBP method Dataset Vanilla ADDT $\ell_{\infty}$
Clean $\ell_{\infty}$ $\ell_{2}$ Clean $\ell_{\infty}$ $\ell_{2}$ Clean $\ell_{\infty}$ $\ell_{2}$
DPDDPM CIFAR-10 85.94 47.27 69.34 85.64 51.46 70.12 84.47 52.64 68.55
@@
Appendix M Additional Experiments under Deterministic White-box Setting
Evaluation across different models
We extend our analysis to include VPSDE and EDM models under the proposed Deterministic White-Box (DW-box) attack setting. The results, presented in Table 22, demonstrate that ADDT consistently improves robustness across different models.
Table 22: DW-box accuracy (%) under $\ell_{\infty}$ perturbations for various models. ADDT consistently improves robustness across all models.
DBP method Vanilla ADDT
DPDDPM 16.80 39.16
DPDDIM 4.98 17.09
DiffPure 22.76 51.63
DPEDM 13.33 32.94
Evaluation across different NFEs
@@
We also investigate the robustness under the Deterministic White-box Setting across varying NFEs. The comparison of performance between vanilla models and ADDT fine-tuned models, shown in Figure 11, highlights that ADDT consistently enhances model performance at different NFEs. This improvement is particularly pronounced at lower NFEs, further confirming that ADDT enables diffusion models to more effectively counter adversarial perturbations.
$10^{1}$$10^{2}$1020304050NFEsAccuracyADDT, White-box (EoT)ADDT, DW-boxVanilla, White-box (EoT)Vanilla, DW-box
@@
Figure 11: Revisiting Deterministic White-box Robustness. ADDT consistently improves robustness under both White-box and Deterministic White-box settings, implying that ADDT strengthens the models’ ability to handle adversarial inputs.
Appendix N Sensitivity Analysis of $\lambda_{unit}$
@@
In Section 6.1 we choose $\lambda_{unit}$=0.03 as the magnitude of such noise is similar to the adversarial perturbation of common adversarial attack setting. We also provide an ablation study here, which shows that the performance of ADDT is insensitive to $\lambda_{unit}$ and gets a consistent improvement.
Table 23: The performance of ADDT fine-tuned DPDDPM with varying $\lambda_{unit}$ values. ADDT shows insensitivity to $\lambda_{unit}$ and consistently improves robust accuracy (%).
NFEs $\ell_{\infty}$ $\ell_{2}$
Vanilla 0.02 0.03 0.04 Vanilla 0.02 0.03 0.04
5 21.78 24.02 30.27 31.25 36.13 41.99 41.99 49.02
10 36.72 40.92 43.07 44.92 55.47 61.72 62.97 64.45
20 45.21 48.14 48.44 50.68 65.23 67.48 69.82 69.24
50 46.78 48.83 50.20 51.07 68.85 69.82 69.04 69.53
100* 47.27 48.93 51.46 50.88 69.34 70.31 70.12 69.92
Appendix O Computational Cost Analysis for Training and Inference
@@
Fine-tuning DDPM and DDIM models using ADDT to achieve near-optimal performance requires 50 epochs and approximately 12 hours of training on 4 NVIDIA GeForce RTX 2080 Ti GPUs. This efficiency matches that of traditional adversarial training approaches and is notably faster than recent adversarial training techniques that utilize diffusion models for dataset augmentation (Wang et al., 2023). However, testing DPDDPM and DPDDIM involves significant computational expense due to the use of Expectation over Transformation (EoT). For instance, validating 1,024 images on the CIFAR10/CIFAR100 datasets takes approximately 5 hours on the same GPU configuration.
One of the key advantages of ADDT is its “train-once” methodology. Once the initial training is complete, ADDT can protect multiple classifiers without requiring additional fine-tuning, as demonstrated in Section 6.1. This is in stark contrast to adversarial classifier training, where each classifier demands individual training.
During inference, models trained with ADDT have a similar complexity to standard DBP. However, their performance gains in accelerated scenarios offer the potential for a reduction in computational overhead. As shown in Section 6.1, DPDDPM + ADDT achieves comparable performance to DPDDPM while requiring only 20 NFEs, resulting in up to an 80% reduction in computation time compared to the 100 NFEs required for DPDDPM.
@@
The code was developed independently by two individuals and mutually verified, with consistent results achieved through independent training and testing. The open-source code is available at https://github.com/LYMDLUT/ADDT.